Calculate the area of the triangle with the following vertices:

(-3, -3), (0, -3), (1, 1)
4 sq units
12 sq units
6 sq units
15 sq units
9 sq units
2 sq units
This is multiply choice, and the possible answers are above as you can see. I have tried everything I can think of to come across the correct answer, but im so confused. I would really apperciate someone explaining how to get the correct answer and what the answer is. Thank you so much!
: Calculate the area of the triangle with the following vertices:

(-3, -3), (0, -3), (1, 1)
4 sq units
12 sq units
6 sq units
15 sq units
9 sq units
2 sq units
This is multiply choice, and the possible answers are above as you can see. I have tried everything I can think of to come across the correct answer, but im so confused. I would really apperciate someone explaining how to get the correct answer and what the answer is. Thank you so much!

Answer by Earlsdon(3816)

When you properly graph these three points and connect them with three straight lines, you will see that the resulting triangle has a base (b) whose length is 3 units and a height (h) of 4 units.
The area of any triangle is given by:
A = (1/2)b*h Substitute b = 3 and h = 4.
A = (1/2)(3)(4) Evaluate.
A = 6sq. units.